Solve the following systems of equations using Cramer's rule:
2x+y=11, x+3y=18.
This is what I came up with:
D=7, Dx=51, Dy=47
with the solution being [7.29, 6.71]
Not sure if this is correct, just need a little help.

Answer by nerdybill(7384)

Your two equations
2x + y = 11
x +3y = 18
.
Your extended matrix:
2 1 11
1 3 18
.
coefficient matrix:
2 1
1 3
Determinant coef (detc): (2)(3)-(1)(1) = 6-1 = 5
.
x matrix
11 1
18 3
Determinant x (detx): (11)(3)-(18)(1) = 33-18 = 15
x = detx/detc = 15/5 = 3
.
y matrix
2 11
1 18
Determinant y (dety): (2)(18) - (1)(11) = 36-11 = 25
y = dety/detc = 25/5 = 5
.
solution:
(x,y) = (3, 5)